OpenStudy (anonymous):

HELP HELP HELP! The sum of two numbers is 12. When a number is subtracted from three times the other, the result is 5. Find the numbers.

OpenStudy (alfie):

a+b = 12 a - 3b = 5.

OpenStudy (alfie):

my bad. 3b- a = 5

OpenStudy (anonymous):

No prob.. then?

OpenStudy (alfie):

then you solve the system.

OpenStudy (anonymous):

\[\frac{31}{4}, \frac{17}{4}\]

OpenStudy (anonymous):

same as b4

OpenStudy (anonymous):

Can you explain clearly? -Alfie?

OpenStudy (anonymous):

Sheg, I need to explain kk?

OpenStudy (anonymous):

two numbers let it be x and y

OpenStudy (anonymous):

sum of these two terms is 12 x + y = 12

OpenStudy (anonymous):

write it once not separately . :(

OpenStudy (anonymous):

X+ Y = 12 3Y -x = 5 add both the equations 4Y = 17 Y = 17 / 4 X = 12 - Y = 12 - 17/4 = 31/4

OpenStudy (alfie):

a+b = 12 3b -a = 5 a = 12-b 3b -12+b = 5 => 4b = 17 => b = 17/4 a = 12-17/4 = 31/4

OpenStudy (anonymous):

OHH! thanks guys. :D